# Building Better: Avoiding Pitfalls in Developing Language Resources when Data is Scarce

### v3

@@ Abstract
Language is a symbolic capital that affects people’s lives in many ways Bourdieu (1977, 1991). It is a powerful tool that accounts for identities, cultures, traditions, and societies in general. Hence, data in a given language should be viewed as more than a collection of tokens. Good data collection and labeling practices are key to building more human-centered and socially aware technologies. While there has been a rising interest in mid- to low-resource languages within the NLP community, work in this space has to overcome unique challenges such as data scarcity and access to suitable annotators. In this paper, we collect feedback from those directly involved in and impacted by NLP artefacts for mid- to low-resource languages. We conduct a quantitative and qualitative analysis of the responses and highlight the main issues related to (1) data quality such as linguistic and cultural data suitability; and (2) the ethics of common annotation practices such as the misuse of online community services. Based on these findings, we make several recommendations for the creation of high-quality language artefacts that reflect the cultural milieu of its speakers, while simultaneously respecting the dignity and labor of data workers.
@@ 1 Introduction
Figure 1: The main themes and targets of our survey. It is designed for NLP researchers and practitioners who have worked on non-high-resource languages (data curation, annotation, and/or model construction). Some of the questions focus on the perspectives of the subset highlighted in the figure, i.e., speakers who focus on their own languages.
@@
There has been an increasing interest in improving the current scope of NLP research with more human-centered design choices Kotnis et al. (2022) and the inclusion of social awareness Yang et al. (2024). As language technologies depend on the quality of the data Hirschberg and Manning (2015) and their alignment with the needs of the speakers, researchers, and other users, the perspectives of these different stakeholders are key to high-quality tools and resources. That is, data selection, annotation, and design choices are traditionally made by the researchers who develop the different artefacts. However, the involvement of those whose native languages are in question is paramount to better design practices Bird and Yibarbuk (2024) as language is part of their culture and identity Bourdieu (1991). That said, when dealing with mid- to low-resource languages in NLP, researchers often make use of the datasets available without necessarily looking into their adequacy, mainly due to resource scarcity. Although progress in NLP for English and other high-resource languages has led to improving standards for corpora quality control and research practices Gebru et al. (2021); Bender and Friedman (2018); Mohammad (2022), one cannot claim the same about the data sources and prevailing practices for mid- to low-resource languages given the current research scope in the field Joshi et al. (2020). Therefore, the NLP artefacts developed for low-resource languages and underrepresented cultures often suffer from a lack of social considerations and over-generalisations due to the over-reliance on data and tools that fail to incorporate the predominant linguistic and cultural features of a given language Bender and Friedman (2018), which may hinder critical progress.
This can further lead to inequality Blasi et al. (2022); Held et al. (2023), sub-optimal experiences with language technologies, and could reinforce a legacy of language hierarchy Kahane (1986).
@@
In this position paper, we shed light on the current limitations of NLP research for mid- to low-resource languages in terms of appropriate data collection, ethical annotation practices, and overall data quality. We reached out to the NLP community involved in NLP projects on under-served languages and conducted a survey to report on the common incentives, limitations, applied norms, and practices (see Figure 1). We outline the survey and present its results. Finally, based on the survey responses, we provide a set of recommendations that focus on (1) fairness and centering of the speakers of the language, (2) choosing suitable data sources, (3) setting fair and realistic expectations when recruiting annotators, and (4) avoiding cultural misrepresentation.
@@ 2 Related Work
Work on ethical practices in AI, ML, and NLP research covers a variety of topic, such as artefact documentation Bender (2011); Bender and Friedman (2018); Gebru et al. (2021); Rogers et al. (2021); Mohammad (2022) and recommendations for best practices Hollenstein et al. (2020); Mohammad (2023). Those that focus on low-resource languages are centered on the general state of NLP research in the area Held et al. (2023); Joshi et al. (2020); Blasi et al. (2022); Doğruöz and Sitaram (2022), limitations in specific tasks such as machine translation Mager et al. (2023), or on the essential question of including people whose languages are in question Mager et al. (2023); Bird (2020, 2022); Bird and Yibarbuk (2024); Lent et al. (2022). Such work sheds light on the peculiarities of low-resource languages with the majority being vernacular languages rather than institutionalised or written Bird and Yibarbuk (2024); Bird (2024). They further advocate for language communities to take over their languagesSchwartz (2022); Markl et al. (2024). For instance, Bird and Yibarbuk (2024) focus on how experts (e.g., linguists, computer scientists) interact with the language communities using participatory design approaches Winschiers-Theophilus et al. (2010), and Cooper et al. (2024) provide recommendations on how to engage with indigenous communities without merely focusing on accuracy. Doğruöz and Sitaram (2022) further point out the importance of not treating language technologies for low-resource languages as scaled-down versions of high-resource ones, and Adebara and Abdul-Mageed (2022) make similar claims with a focus on features that are specific to African languages.
In addition to the language speakers, other work focuses on users such as Blaschke et al. (2024) who focus on dialect speakers and emphasise the need to involve end users in designing language technologies. Moreover, Yang et al. (2024) define social awareness and advocate for refraining from treating language in NLP as a computational problem only. In this paper, we strengthen the above discussion by shifting the focus to the practical challenges faced by NLP researchers and practitioners working on mid- to low-resource languages by borrowing practices from social science Cetina (1999) to study the methodological practices and issues in the field. To the best of our knowledge, there is limited work investigating low-resource research while trying to connect to online communities, except for three case studies discussed by Birhane et al. (2022), and work by Lent et al. (2022), who analyse 38 responses collected on Facebook and Twitter/X. By analysing this feedback, we aim to present practical recommendations that emphasise transparency and ethically grounded practices for building more human-centered language technologies.
@@ 3 Survey
Our main goal is to investigate the current issues and problematic practices in NLP research for mid- to low-resource languages and provide potential solutions. Therefore, we reached out to the NLP community from June to October 2024 on X, LinkedIn, Google groups and Slack channels of NLP communities, and by direct emails. We targeted researchers working on mid- and low-resource languages, language variants, dialects, and vernaculars and surveyed how research is conducted. Participants report on common practices, incentives, and issues that stand out. Then, we present a quantitative and qualitative analysis of the responses.
@@ 3.1 Respondents
Respondents are NLP researchers and practitioners involved in the data collection, annotation, model construction, or other research questions related to mid- to low-resource languages. Some may have additionally also conducted research for high-resource languages. Note that the respondents may or may not speak the languages.
@@ 3.2 Survey Structure
We ask the respondents about (1) their previous experiences in the area, (2) current problems and limitations relevant to their language(s) of interest, (3) the motivation behind their involvement in various projects, and (4) how they were credited for tasks that are often specific to low-resource languages, e.g., compensation for annotations done via online community forums. Note that the respondents decide what a mid- to low-resource language means to them.
@@ 3.2.1 General Questions
Respondents were asked to optionally fill in their names and contact information for a potential follow-up. Then, they were asked about:
@@
  • whether they are/were part of any online community,
  • whether the projects that they worked on are from industry, academia, or both,
  • the kind of NLP tools that are or would be relevant and useful in their language(s) on interest,
@@ 3.2.2 Reporting on Incentives and Potential Limitations
As we are aware of potential limitations in NLP for mid- to low-resource languages Blasi et al. (2022), we examined whether the respondents work in the area due to any limitations observed in available NLP tools or conducted research in their language(s) of interest. Note that these questions were optional as researchers may work on mid- to low-resource languages for various other reasons. We asked the participants to report on:
  • any observed limitations and optionally list some tools or resources on language(s) of interest as examples,
@@ 3.2.3 Reporting on Credit Attribution
We asked the respondents about how often they were properly credited for their work. Further, as reaching out to online communities is common to projects that include mid- to low-resource languages, we asked whether the participants were involved in past projects through online community platforms (for data collection, annotation, model construction, etc.). This is because involving communities in NLP and ML projects is relatively new to the field and can therefore be abused as there are no clear standards regarding data workers in such contexts. Therefore, we ask the respondents about:
  • How often did they receive credit for their contributions? E.g., whether they received proper financial compensation for annotating a dataset.
@@
  • How long did the process take especially when they were not properly compensated?
@@ 4 Findings
Table 1: Reported project affiliations, tasks in which the annotators were involved, and the different motivations or incentives. Note that percentages do not sum up to one as respondents could report on more than one project.
We received 81 responses from researchers working on a wide range of mid- to low-resource languages and language families. Even though including contact information was optional, more than 90% of the respondents chose not to reply anonymously, and 80% asked for updates on the project. Table 1 shows the distribution of responses to questions on project affiliations, the tasks in which the respondents were involved, and their motivations for working on mid- to low-resource languages. Note that percentages do not sum up to one as respondents could report on more than one project.
  1. 1.
    the tasks that they work on (i.e., data creation 47%, collection 33%, annotation 33%, and model construction 9%);
  2. 2.
    the reasons why they chose to work on their language(s) of interest (i.e., 81% scientific interest or curiosity, 72% for building language technologies, 60% because of observed limitations in resources focusing on their language(s) of interest, and 59% for LLM research).
Figure 2: The main locations where the languages that our survey respondents work on are spoken.
@@ 4.1.1 Projects
The respondents could report on one or many projects they have been involved in. As shown in Table 1, Most respondents have worked on academic projects, with a third on collaborations between industry and academia or both types of projects.  [^1] Note that although ¿50% of the respondents named the projects they participated in and did not mind sharing their names publicly, we do not disclose this information to protect the anonymity of our respondents.
@@ 4.1.2 Languages
Among the 81 responses, the respondents worked on >70 low-resource languages they specifically named (see Appendix). Figure 2 illustrates the main locations where these languages are spoken. The languages include variants, dialects, and vernaculars (e.g., country-specific Arabic dialects), truly low-resource languages (e.g., Welsh, Yoreme Nokki, Setswana), and mid to low-resource ones (e.g., Amharic, Indonesian). In addition, about 12% of the respondents reported working on language families and language branches such as South Asian languages, all Gaeilige dialects, or Arabic/English variations. A high percentage of the respondents work on high-resource languages as well (e.g., English, French, Spanish, and Modern Standard Arabic).
@@ 4.2 Incentives and Potential Limitations
When asking the respondents about why they work on NLP for mid to low-resource languages, we provide them with a checklist from which they could choose more than one option and add their own entry. We report on the frequent motivations and practices that are only adopted in non-high-resource settings due to, e.g., data scarcity. We identify problematic instances and analyse the possible reasons behind some.
Figure 3: Frequency of each incentive that was found in our survey responses. Note that the percentages do not sum up to 100 as the respondents could choose more than one option.
When further analysing the common motivations, we report more detailed numbers in Figure 3. Among those who were motivated by scientific curiosity or interest in Table 1 there were those whose interest was in NLP/CL/ML research (68%) and those whose interest was in languages (68%). Note that the two are not mutually exclusive.
Moreover, among the respondents who are interested in building language technologies, most of them were more interested in building technologies for their own language(s) (60%) as opposed to building technologies for as many languages as possible (38%). This is particularly interesting as it constitutes evidence of the power of language as a symbolic capital Bourdieu (1991), which can sometimes manifest in the feeling of “a duty” that one has towards their language. Other frequent motivations include marked limitations in language resources and tools in the language(s) of interest (60%) and the willingness to contribute to research on LLMs (59%).
@@ 4.2.1 Reported Limitations
More than 60% of the respondents reported working on low-resource languages due to marked limitations in currently available resources for their language(s) of interest. To shed light on these limitations, we showed the respondents a predefined list of possible shortcomings as well as a text box where they could add any observed limitations. As shown in Figure 4: the predominant limitation is data scarcity (78%). This is followed by the lack of representativeness of the data (58%), the under-performance of the available tools (54%), their misalignment with the users’ needs (54%), the low quality of the annotations (25%), and the lack of the usefulness of the data (18%).
@@ 4.2.2 Qualitative Analysis of the Limitations
We provided the respondents with free text sections where they could report examples of tools or resources that suffer from the limitations that they mentioned to justify their choices. When manually processing the answers, we noticed the following themes:
@@
    • the reliance on machine translation tools and LLMs to build resources for under-resourced languages;
@@
    • the potential misuse of online community services.
@@
    Fundamental problems related to NLP research on mid to low-resource languages: such as the lack of funding often due to the “low prestige” language dilemma—the false notion that some languages or language varieties are more important than others.
@@
As many languages are not institutional but rather vernacular Bird and Yibarbuk (2024), data collection presents considerable challenges when solely relying on textual data for low-resource languages, e.g., Bantu languages.
Further, the focus on English and the reliance on translated data harms the quality of the generated datasets as they do not capture the subtle peculiarities of a given language. Another issue is what is commonly called “the curse of multilinguality" as the commonly used multilingual tools do not perform as well as the monolingual ones.
@@
Lack of representativeness and naturalness as well as “attention to details" were commonly reported in the responses. The respondents reported a lack of awareness of language variants and cultural aspects when building a language-specific artefact; the reliance on the standardised version of a given language due to power dynamics (more power in the hands of well-funded institutions and established researchers); the presence of offensive utterances in the data due to a lack of data filtering; and potentially wrong assumptions about a language or a culture. Further, the time-specific context and usage of some languages, such as ancestral ones (e.g., Coptic), have considerably changed and one has to take these facts into account. In addition, datasets may be collected from inadequate sources or could be aligned with Western values, standards, or expectations. This can be due to the previous points as well as power differentials or even well-intended educated guesses without a careful or deeper examination carried along with locals and native speakers.Finally, researchers rely on personal connections as it is hard to impossible to find native speakers of mid to low-resource languages on commonly used annotation platforms such as Amazon Mechanical Turk, Polific, and others. Added to this reason, the lack of funding leads researchers to turn to online community work. This practice has been at the center of major NLP contributions in recent years. However, despite its clear benefits, the absence of well-established standards puts community members at risk as their efforts may not be properly recognised.
@@
Many respondents reported that conducting research in mid- to low-resource languages often entailed high costs of data curation, potential reach out to local communities, the need for resources, and the cost of the datasets that are not freely available.

### v5

@@ Abstract
Language is a symbolic capital that affects people’s lives in many ways Bourdieu (1977, 1991). It is a powerful communication means that accounts for identities, cultures, traditions, and societies in general. Hence, data in a given language should be viewed as more than a collection of tokens. Good data collection and labeling practices are key to building more human-centered and socially aware technologies. While there has been a rising interest in under-resourced languages within the NLP community, work in this space has to overcome unique challenges such as data scarcity and access to suitable annotators. In this paper, we collect feedback from those directly involved in and impacted by NLP artefacts for medium and low-resource languages. We conduct a quantitative and qualitative analysis of the responses and highlight the main issues related to (1) data quality such as linguistic and cultural data suitability; and (2) the ethics of common annotation practices such as the misuse of online community services. Based on these findings, we make several recommendations for the creation of high-quality language artefacts that reflect the cultural milieu of its speakers, while simultaneously respecting the dignity and labor of data workers.
@@ 1 Introduction
Figure 1: The main stakeholders and their role in a data annotation project. We reached out to the CL community in our survey. It is designed for NLP researchers and practitioners who have worked on low- to medium-resource languages. Some of the questions focus on the perspectives of the subset highlighted in the figure, i.e., speakers who focus on their own languages.
@@
Over the past few years, there has been an increasing interest in improving the scope of NLP research with more human-centered design choices Kotnis et al. (2022) and social awareness Yang et al. (2024). As language technologies depend on the quality of the data Hirschberg and Manning (2015) and their alignment with the needs of the speakers, researchers, and other users, the perspectives of these different stakeholders are key to high-quality tools and resources. That is, data selection, collection, annotation, and design choices are traditionally made by the researchers who develop the different artefacts. However, the involvement of those whose native languages are in question is paramount to better design practices Bird and Yibarbuk (2024) since language is part of their culture and identity Bourdieu (1991). That said, when dealing with low to mid-resource languages in NLP, researchers often make use of the datasets available without necessarily looking into their quality or questioning their suitability, mainly due to resource scarcity. Although progress in NLP for English and other high-resource languages has led to improving standards for corpora quality control and research practices Gebru et al. (2021); Bender and Friedman (2018); Mohammad (2022), one cannot claim the same about the data sources and prevailing practices for low- to mid-resource languages given the current research scope in the field Joshi et al. (2020). Therefore, the NLP artefacts developed for under-resourced languages and underrepresented cultures often suffer from a lack of social considerations and cause over-generalisations due to the over-reliance on data and tools that fail to incorporate the predominant linguistic and cultural features of a given language Bender and Friedman (2018), which may hinder critical progress. This can further lead to inequality Blasi et al. (2022); Held et al. (2023), sub-optimal experiences with language technologies, and could reinforce a legacy of language hierarchy Kahane (1986).
@@
In this position paper, we shed light on the current limitations of NLP research for low- to mid-resource languages in terms of appropriate data collection, ethical annotation practices, and overall data quality. We reached out to the NLP community involved in NLP projects on under-served languages and conducted a survey to report on the common incentives, limitations, applied norms, and practices (see Figure 1). We outline the survey and present its results. Finally, based on the survey responses, we provide a set of recommendations that focus on (1) fairness and centering of the speakers of the language, (2) choosing suitable data sources, (3) setting fair and realistic expectations when recruiting annotators, and (4) avoiding cultural misrepresentation.
@@ 2 Related Work
Work on ethical practices in AI, ML, and NLP research covers a variety of topics, such as artefact documentation Bender (2011); Bender and Friedman (2018); Gebru et al. (2021); Rogers et al. (2021); Mohammad (2022) and recommendations for best practices Hollenstein et al. (2020); Mohammad (2023). Those that focus on low-resource languages are centered on the general state of NLP research in the area Held et al. (2023); Joshi et al. (2020); Blasi et al. (2022); Doğruöz and Sitaram (2022), limitations in specific tasks such as machine translation Mager et al. (2023), LLM research Mihalcea et al. (2024), or on the essential question of including people whose languages are in question Mager et al. (2023); Bird (2020, 2022); Bird and Yibarbuk (2024); Lent et al. (2022). Such work sheds light on the peculiarities of low-resource languages with the majority being vernacular languages rather than institutionalised or written Bird and Yibarbuk (2024); Bird (2024). They further advocate for language communities to take over their languages Schwartz (2022); Markl et al. (2024); Mihalcea et al. (2024). For instance, Bird and Yibarbuk (2024) focus on how experts (e.g., linguists, computer scientists) interact with the language communities using participatory design approaches Winschiers-Theophilus et al. (2010), and Cooper et al. (2024) provide recommendations on how to engage with indigenous communities without merely focusing on accuracy. Doğruöz and Sitaram (2022) further point out the importance of not treating language technologies for low-resource languages as scaled-down versions of high-resource ones and account for their specificity. Similarly, Adebara and Abdul-Mageed (2022) make related claims with a focus on features that are peculiar to African languages (e.g., tones). In addition to the language speakers, other work focuses on users such as Blaschke et al. (2024) who highlight the needs of dialect speakers and the importance of involving them in designing language technologies. Moreover, Yang et al. (2024) define social awareness and advocate for refraining from treating language in NLP as a computational problem only. In this paper, we strengthen the above discussion by shifting the focus to the practical challenges faced by NLP researchers and practitioners working on low and mid-resource languages by borrowing practices from social science Cetina (1999) to study the methodological practices and issues in the field. To the best of our knowledge, there is limited work investigating NLP research for low-resource languages while trying to connect to online communities, except for three case studies of participatory frameworks (e.g., Masakhane), and work by Lent et al. (2022), who analyse 38 responses by Creole speakers on their experiences with language technologies. By analysing the respondents’ feedback, we present practical recommendations that emphasise transparency and ethically grounded practices for building more human-centered NLP artefacts for under-served languages.
@@ 3 Survey
Our main goal is to investigate the current issues and problematic practices in NLP research for low- to mid-resource languages and provide potential solutions. Therefore, we reached out to the NLP community (i.e., *CL networks) from June to October 2024 on X, LinkedIn, Google groups, and Slack channels of online NLP communities, and by direct emails. We targeted researchers and annotators working on low- and mid-resource languages, language variants, dialects, and vernaculars, and surveyed how research is conducted. Participants report on common practices, incentives, and issues that stand out. Then, we present a quantitative and qualitative analysis of the responses.
@@ 3.1 Respondents
Respondents are NLP researchers and practitioners involved in the data collection, annotation, model construction, or other research questions related to under-served languages.
@@ 3.2 Survey Structure
We ask the respondents about (1) their previous experiences in the area, (2) current problems and limitations related to their language(s) of interest, (3) the motivation behind their involvement in various projects, and (4) how they were credited for tasks that are often specific to low-resource languages, e.g., compensation for annotations done via online community forums or participatory frameworks.
Note that we left it to the respondents to decide on what represents a mid- to low-resource language since there is no standard definition. E.g., most researchers would agree that Tamasheq is a low-resource language, while they may not have the same opinion about Malaysian.
@@ 3.2.1 General Questions
Respondents could optionally fill in their names and contact information for a potential follow-up. Then, they were asked about:
@@
  • whether they are/were part of any online community (i.e., participatory research frameworks),
  • whether the project(s) they worked on are from industry, academia, or both,
  • the kind of NLP tools that are or would be relevant and useful in their language(s) of interest,
@@ 3.2.2 Reporting on Incentives and Potential Limitations
As we are aware of potential drawbacks in NLP for mid- to low-resource languages Blasi et al. (2022), we examined whether the respondents work in the area due to any limitations observed in available NLP tools in their language(s) of interest. Note that these questions were optional as researchers may work on low- to mid-resource languages for various other reasons. We asked the participants to report on:
  • any observed limitations and optionally list some tools or resources in their language(s) of interest as examples,
@@ 3.2.3 Reporting on Credit Attribution
We asked the respondents about how often they were properly credited for their work. Further, as reaching out to online participatory frameworks is common to projects that focus on under-resourced languages, we asked whether the participants were involved in past projects through participatory frameworks. This is because involving communities in NLP and ML projects is relatively new to the field and can therefore be abused (purposefully or not) due to the lack of clear standards regarding data workers in such contexts. Therefore, our questions were the following:
  • How often did the respondents receive credit for their contributions? E.g., financial compensation for annotating data.
@@
  • How long did the process take especially when they were not properly credited?
@@ 4 Findings
Table 1: Reported project affiliations, tasks in which the annotators were involved, and the different motivations or incentives. Note that percentages do not sum up to one as respondents could report on more than one project in both industry and academia.
We received 81 responses from researchers working on a wide range of languages and language families. Even though including contact information was optional, more than 90% of the respondents chose not to reply anonymously, and 80% asked for updates on the project. Table 1 shows the distribution of responses to questions on project affiliations, the tasks in which the respondents were involved, and their motivations for working on under-resourced languages. Note that percentages do not sum up to one as respondents could report on more than one project. That is, participants could be involved in several tasks and projects. As shown in Table 1, most participants were involved in dataset curation mainly motivated by scientific interest or curiosity, and for building language technologies because of observed limitations in resources dedicated to their language(s) of interest.
Figure 2: The main locations where the languages of interest of our respondents are spoken.
@@ 4.1.1 Projects
The respondents could report on one or many projects they have been involved in. As shown in Table 1, Most respondents have worked on academic projects, with a third on collaborations between industry and academia or both types of projects. [^1] Note that although ¿50% of the respondents named the projects they participated in, we do not disclose it to protect the anonymity of our respondents.
@@ 4.1.2 Languages
Among the 81 responses, the respondents worked on >70 low-resource languages they specifically named (see Appendix). Figure 2 illustrates the main locations where these languages are spoken. These include variants, dialects, and vernaculars (e.g., country-specific Arabic dialects), truly low-resource languages (e.g., Welsh, Yoreme Nokki, Setswana), and low- to medium-resource ones (e.g., Amharic, Indonesian). In addition, about 12% of the respondents reported working on language families and language branches such as South Asian languages, all Gaeilige dialects, or Arabic/English variations. A high percentage of the respondents work on high-resource languages in parallel.
@@ 4.2 Incentives and Potential Limitations
When asking the respondents about why they work on NLP for under-resourced languages, we provide them with a checklist from which they could choose more than one option or add their own entry. We report on the frequent motivations and practices that are only adopted in low- to mid-resource settings due to, e.g., data scarcity. We identify problematic instances and analyse the possible reasons behind some.
Figure 3: Frequency of each incentive. Note that the percentages do not sum up to 100 as the respondents could choose more than one option.
When further examining the common motivations, we report more detailed numbers in Figure 3. Among those who were motivated by scientific curiosity or interest in Table 1 there were those whose interest was in NLP/CL/ML research (68%) and those whose interest was in languages (68%). Note that the two are not mutually exclusive.
Moreover, for the respondents whose motivation was building language technologies, most of them were more interested in building technologies for their own language(s) (60%) as opposed to building technologies for as many languages as possible (38%). This is particularly interesting as it constitutes evidence of the power of language as a symbolic capital Bourdieu (1991), which can sometimes manifest in the feeling of “a duty” that one has towards their language. Other frequent motivations include marked limitations in language resources and tools in the language(s) of interest (60%) and the willingness to contribute to LLM research (59%).
@@ 4.2.1 Reported Limitations
More than 60% of the respondents reported working on low-resource languages due to marked limitations in currently available resources for their language(s) of interest. To shed light on these limitations, we showed the respondents a predefined list of possible shortcomings as well as a text box where they could add any observed limitations. As shown in Figure 4: the predominant limitation is data scarcity (78%); followed by the lack of representativeness of the data (58%) that can manifest in, e.g., unnatural or translationese data instances; the under-performance of the available tools (54%); their misalignment with the users’ needs (54%); the low quality of the annotations (25%); and the lack of the usefulness of the data (18%).
@@ 4.2.2 Qualitative Analysis of the Limitations
Respondents on getting credit for projects they were involved in. We provided the respondents with free text sections where they could report examples of tools or resources that suffer from the limitations that they mentioned to justify their choices. When manually processing the answers, we noticed the following recurring topics:
@@
    • the reliance on machine translation tools and LLMs to build resources;
@@
    • the potential misuse of online communities and participatory framework services.
@@
    Fundamental problems related to NLP research on low- to mid-resource languages: such as the lack of funding often due to the “low prestige” language dilemma—the false notion that some languages or language varieties are more important than others. Interestingly, this was equally observed in projects from academia and industry.
@@
Since many under-resourced languages are not institutional but rather vernacular Bird and Yibarbuk (2024), collecting data presents considerable challenges when one solely relies on textual data for, e.g., Bantu languages. Further, the focus on English and the reliance on translated data harms the quality of the generated datasets as they do not capture the subtle peculiarities of a given language. It is important to note that what is translated and whether it was further verified by a fluent speaker makes a difference as translating Wikipedia texts can be easier than translating conversational, informal, or religious texts Hutchinson (2024).
@@
Lack of representativeness and unnaturalness of the data were commonly reported in the responses. The respondents reported a lack of awareness of language variants and cultural aspects when building a language-specific artefact; the reliance on the standardised version of a given language due to power dynamics (more power in the hands of well-funded institutions and established researchers); the presence of offensive utterances in the data due to a lack of data filtering; and potentially wrong assumptions about a language or a culture. For instance, the time-specific context and usage of some languages, such as ancestral ones (e.g., Coptic), have considerably changed and one has to take these facts into account. In addition, datasets may be collected from inadequate sources or could be aligned with Western values, standards, or expectations. This can be due to power differentials or a lack of deeper examination carried along with locals and native speakers. Finally, researchers rely on personal connections as it is hard to impossible to find fluent speakers of low- to mid-resource languages on commonly used annotation platforms such as Amazon Mechanical Turk and Polific. Added to this reason, the lack of funding leads researchers to turn to participatory frameworks. This practice has been at the center of major NLP contributions in recent years Birhane et al. (2022). However, despite its benefits for people with common research interests, the absence of well-established standards puts vulnerable community members at risk as their efforts may not be properly recognised.
@@
Many respondents reported that conducting research in low- to mid-resource languages often entails high costs of data curation and potential outreach to local communities. Further, when resources for an under-resourced language exist, they are often not freely available.